OpenStudy (anonymous):

please help will give medals how many solutions does the system of equations have 3x = -12y + 15 and x + 4y = 5 how many solutions does the system of equations have y = 6x + 2 and 3y - 18x = 12 how many solutions does the system of equations havex - 4y = 12 and 5x - 20y = 60 how many solutions does the system of equations have y - 6x = -3 and 4y - 24x = - 16

OpenStudy (texaschic101):

3x = - 12y + 15 3x + 12y = 15 reduces to x + 4y = 5 this matches the other equation, there, they are on the same line and have INFINITE SOLUTIONS ============ y = 6x + 2 -6x + y = 2 6x - y = -2 3y - 18x = 12 -18x + 3y = 12 18x - 3y = -12 reduces to 6x - y = -4 The equations do not match, therefore, there is NO SOLUTION ============ x - 4y = 12 5x - 20y = 60 reduces to x - 4y = 12 same equation, same line, INFINITE SOLUTIONS =========== y - 6x = -3 -6x + y = -3 6x - y = 3 4y - 24x = -16 -24x + 4y = -16 24x - 4y = 16 reduces to 6x - y = 4 NO SOLUTIONS